Cost of Units Completed and in Process

The charges to Work in Process—Assembly Department for a period, together with information concerning production, are as follows. All direct materials are placed in process at the beginning of production.

Work in Process—Assembly Department
Bal., 5,000 units, 40% completed 11,900 To Finished Goods, 115,000 units ?
Direct materials, 118,000 units @ $1.5 177,000
Direct labor 200,100
Factory overhead 77,820
Bal., ? units, 35% completed ?

Cost per equivalent units of $1.50 for Direct Materials and $2.40 for Conversion Costs.

a. Based on the above data, determine the different costs listed below.

If required, round your interim calculations to two decimal places.

1. Cost of beginning work in process inventory completed this period $
2. Cost of units transferred to finished goods during the period $
3. Cost of ending work in process inventory $
4. Cost per unit of the completed beginning work in process inventory (Rounded to the nearest cent.) $

b. Did the production costs change from the preceding period?

c. Assuming that the direct materials cost per unit did not change from the preceding period, did the conversion costs per equivalent unit increase, decrease, or remain the same for the current period?
Answer #1

Answer:

Note : Using FIFO Method

Balance Closing Unit = Opening + Introduced - Closing = 5,000 + 118,000 - 115,000 = 8,000

Cost per Unit

Particulars Units %age completed during the year Material %age completed during the year Labor + Factory overhead
Cost 177,000 277,920
Equivalent units
Opening 5,000 0 0 60% 3,000
Introduced 110,000 100% 110,000 100% 110,000
Closing 8,000 100% 8,000 40% 3,200
Total 123,000 118,000 116,200
Cost per unit 1.50 2.40

a)

i) Cost of beginning work in progress

= opening cost + opening equivaent material units X Cost per unit + Opening equivalent overhead unit X cost per unit

= 11,900 + ( 0 X 1.50 ) + ( 3,000 X 2.40)

= 19,100

ii) Cost of units transfered to finished goods i.e 115,000 units

= opening completed goods cost + cost of Units started and completed during year

= 19,100 + (110,000 X 1.50) + ( 110,000 X 2.40)

= 448,100

iii) Cost of ending work in progress

= Cost of equivalent material unit X cost per unit + Cost of equivalent overhead units X Cost per unit

= 8,000 X 1.50 + 3,200 X 2.40

= 19,680

IV) Cost per unit of the completed beginning work in process inventory

= Opening Cost / Opening completed units

= 19,100 / 5,000

= 3.82

b) Yes it is changed by ( 1.50 + 2.40) - 3.82= $ 0.08

C) Increases

Previous cost = 3.82 - 1.50 = 2.32

Current cost = 2.32

increase = 0.08